OpenStudy (ammarah):

Describe the end behavior of the graph of the polynomial function by completing these statements: f(x)--------->____as x------>-x and f(x)----->______as x----> +x 1. f(x) =-5x^3 2. f(x)=2x^5-7x^2-4x

OpenStudy (ammarah):

@Easyaspi314

OpenStudy (ammarah):

@ganeshie8

OpenStudy (mathstudent55):

Do you mean x ---> \(-\infty\), and x---> \(\infty\)?

OpenStudy (ammarah):

yess

OpenStudy (ammarah):

do u get this i dont get it at all i have like 5 questions like this

OpenStudy (mathstudent55):

Let's look at 1.

OpenStudy (mathstudent55):

f(x) = -5x^3 Just look at x^3. As x gets larger and larger, what happens to x^3?

OpenStudy (ammarah):

it gets smaller???? how would u know?

OpenStudy (mathstudent55):

Smaller? Try it. If x = 1, x^3 = 1 If x = 10, x^3 = 1000 If x = 100, x^3 = 1000000 So as x gets larger, x^3 gets larger too.

OpenStudy (ammarah):

ohh okay so it would be the positive sign of infenity in the first blank and negative in the second

OpenStudy (mathstudent55):

No. We're not done yet.

OpenStudy (mathstudent55):

In 1., you have f(x) = -5x^3 We are looking only at x^3 so far.

OpenStudy (mathstudent55):

As x gets larger, x^3 gets larger. As x goes to infinity, x^3 also goes to infinity. So far you follow?

OpenStudy (ammarah):

yes

OpenStudy (mathstudent55):

Now remember that the functuion is f(x) = -5x^3. Whatever x is, you need to cube it first, then multiply it by negative 5. That means that as x goes to infinity, x^3 goes to infinity, and -5x^3 goes to negative infinity.

OpenStudy (mathstudent55):

This is the answer to the second part of 1. As x ---> \(+\infty\), f(x) ---> \(-\infty\)

OpenStudy (mathstudent55):

For the first part of 1., let's look again at x and f(x) As x goes to \(- \infty\), x^3 also goes to \(- \infty\), but the function is f(x) = -5x^3. As x^3 goes to \(- \infty\), -5x^3 goes to \( +\infty\), so for the first part of 1. the naswer is f(x) ---> \(+ \infty\) as x ---> \(- \infty\)

OpenStudy (ammarah):

what would be the second question would it be positive infinity and then negative?:???

OpenStudy (mathstudent55):

Let's do the part where x approaches positive infinity first. As x gets larger, the first term of the polynomial, 2x^5, because of the 5th power gets very large. So as x goes to infinity, f(x) also goes to infinity. For the first part, x goes to negative infinity. x^5 also goes to negative infinity, so as x goes to negative infinity, f(x) also approaches negative infinity.

OpenStudy (ammarah):

why negative i dont get it